If the pupil in a class sends a New Year greeting cards to every classmate ,the total number of cards sent out will be 870. Find the number of pupils in the class.
Answer by checkley77(12844)

EVERY PUPIL SENDS CARDS TO (X-1) PUPILS.
X IS THE TOTAL NUMBER OF PUPILS.
X(X-1)=870
X^2-X-870=0
(X-30)(X+29)=0
X-30=0
X=30 NUMBER OF PUPILS.